# Scaling Vision Mamba Across Resolutions via Fractal Traversal

### v1

@@ Abstract
Vision Mamba has recently emerged as a promising alternative to Transformer-based architectures, offering linear complexity in sequence length while maintaining strong modeling capacity. However, its adaptation to visual inputs is hindered by challenges in 2D-to-1D patch serialization and weak scalability across input resolutions. Existing serialization strategies such as raster scanning disrupt local spatial continuity and limit the model’s ability to generalize across scales. In this paper, we propose FractalMamba++, a robust vision backbone that leverages fractal-based patch serialization via Hilbert curves to preserve spatial locality and enable seamless resolution adaptability. To address long-range dependency fading in high-resolution inputs, we further introduce a Cross-State Routing (CSR) mechanism that enhances global context propagation through selective state reuse. Additionally, we propose a Positional-Relation Capture (PRC) module to recover local adjacency disrupted by curve inflection points. Extensive experiments on image classification, semantic segmentation, object detection, and change detection demonstrate that FractalMamba++ consistently outperforms previous Mamba-based backbones, particularly under high-resolution settings.
@@ 1 Introduction
Figure 1: Visualization of correlation between the final state and intermediate state tokens in the deepest feature layer under different input resolutions. The 16×16$\times$16 × and 32×32$\times$32 × maps correspond to the output token grids for input resolutions of $512\times 512$ and $1024\times 1024$, respectively. Without CSR, correlations are biased and localized due to information fading across long sequences. With CSR enabled, strong and distributed correlations are observed, demonstrating that global contextual information is effectively propagated across the sequence. “W/o” means without operation.
@@
Figure 2: Illustration of the structural limitation of fractal curves. Certain corner points along the fractal traversal path (highlighted by purple circles) experience disrupted adjacency relationships. Although spatially close in the 2D image grid, these patches become distant in the serialized sequence, impairing local continuity and information exchange.
@@
  • We conduct a more comprehensive set of experiments than prior work, evaluating FractalMamba++ across multiple parameter scales on diverse visual tasks including classification, segmentation, detection, and change detection.
@@ 3.3 Cross-State Routing
TABLE II: Image classification performance (Top-1 Accuracy) on ImageNet-1k under varying input resolutions. FLOPs are measured at input resolution of $224\times 224$.
Method Publication Param. FLOPs Input Resolution
2242 2562 3842 5122 6402 7682 10242 12802 14082 15362
Mamba® CVPR’25 29M 4.6G 81.1 45.7 25.4 12.8 7.8 5.3 2.8 1.8 1.6 1.4
FractalMamba++ Year’25 30M 4.8G 83.0 83.3 84.1 83.9 83.0 81.9 78.8 74.3 71.3 67.5
FractalMamba++ Year’25 11M 1.6G 79.5 80.6 82.0 81.3 80.1 78.3 73.3 66.3 61.7 56.1
ViM ICML’24 7M 1.5G 76.1 76.3 70.4 67.4 51.4 30.6 16.1 7.2 4.1 1.8
FractalMamba++ Year’25 7M 1.0G 77.3 78.4 79.5 78.4 76.4 73.7 66.5 55.2 48.1 42.5
@@ 3.4 Positional-Relation Capture
where $\operatorname{Re}[\cdot]$ denotes the real part and is the complex conjugate. This formulation introduces a rotation-phase difference proportional to the relative distance $(n-m)$, thereby encoding relative positions directly into token similarity.
@@ 4 Experiments
We conduct a series of experiments to evaluate and compare FractalMamba++ against several established Mamba-based benchmark models across a diverse range of vision tasks. These tasks include image classification, object detection, remote sensing binary change object detection, and semantic segmentation. Each task requires a different input resolution, which provides a comprehensive testbed to evaluate the multi-scale adaptability of our model. In particular, the remote sensing binary change object detection task involves high-resolution input images of size 1024 $\times$ 1024. This setting places strong demands on the model’s ability to maintain spatial consistency and capture long-range structural relationships. All experiments are conducted using 8 NVIDIA H800 GPUs.
@@ 4.1 Image Classification
FractalMamba++ consistently achieves competitive or superior performance across all tested resolutions. At lower resolutions (e.g., $224^{2}$, $384^{2}$), it performs on par with state-of-the-art Mamba-based models. However, as input resolution increases, FractalMamba++ shows significantly slower degradation in accuracy compared to all baselines. For instance, at $1024^{2}$, the 30M variant of FractalMamba++ achieves 78.8% top-1 accuracy, outperforming SpatialMamba (66.1%), GrootV (67.9%), and the original VMamba (62.3%). Similar trends are observed at 11M and 7M scales. This superior scalability highlights the effectiveness of our design. The fractal-based patch serialization preserves spatial locality across scales, while the CSR and PRC modules enhance long-range context propagation and mitigate structural disruptions caused by curve inflections. Together, these modules allow FractalMamba++ to maintain strong performance under extreme-resolution inputs, confirming its multi-scale adaptability and robustness for real-world visual applications.
To complement accuracy evaluations, we further compare the inference efficiency of various Mamba-based models on ImageNet-1K at $224\times 224$ resolution. As shown in Table III, we report frames per second (FPS) on a single H800 GPU with batch size 32. FractalMamba++ achieves competitive inference speeds within each parameter scale, while offering consistently superior multi-resolution performance. In the $\sim$30M group, FractalMamba++ runs at 302 FPS, slower than VMamba (1916 FPS) and MambaVision (3426 FPS), but significantly faster than MILA (47 FPS) and on par with GrootV (315 FPS). In the medium-scale group ($\sim$11M), FractalMamba++ achieves 511 FPS, outperforming ViM (398 FPS) in the small-scale group ($\sim$7M), while Efficient VMamba maintains the highest FPS across all groups. Although FractalMamba++ incurs moderate latency due to fractal serialization and additional modules (CSR, PRC), it still delivers a favorable trade-off between speed and accuracy, especially under high-resolution scenarios where other models degrade sharply. Notably, the current implementation of the CSR module involves custom CUDA operators, which, while functional, have not been fully optimized. This suggests that there remains considerable headroom for further acceleration through kernel-level refinement. This highlights the practical deployment potential of FractalMamba++ in resource-constrained settings.
@@ 4.2 Object Detection
Table IV presents detection and segmentation results for FractalMamba++ and several representative Mamba-based models, including VMamba [26], MSVMamba [34], Efficient VMamba [38], SpatialMamba [35], DefMamba [56], and FractalMamba [32]. Performance is reported in terms of mean average precision (mAP) for bounding boxes and instance masks, including scores at 50 and 75 percent IoU thresholds. FractalMamba++ achieves consistent improvements under both training settings. Under the 1× schedule, the 41M model obtains 47.8 mAP for box prediction and 43.2 mAP for mask prediction, outperforming Spatial Mamba (47.6 and 42.9), MSVMamba (46.9 and 42.2), and VMamba (46.5 and 42.1). Under the 3× schedule, FractalMamba++ further improves to 49.5 for box mAP and 44.1 for mask mAP, achieving the highest performance among comparable models. Similar trends are observed at the 31M scale, where FractalMamba++ achieves 46.8 for box mAP and 42.2 for mask mAP, compared with 46.3 and 41.8 from MSVMamba. These results confirm the effectiveness and generalizability of FractalMamba++ across dense prediction tasks. The fractal-based serialization improves spatial structure retention, while the CSR and PRC modules enable stronger context modeling and structural continuity, which are essential for high-resolution visual understanding.
TABLE V: Semantic segmentation results on ADE20K using UPerNet with input resolution $512\times 2048$. FLOPs are computed under the same setting. SS and MS denote single-scale and multi-scale inference.
@@ 4.3 Semantic Segmentation
Table V reports results across three model capacities. At the Tiny scale (62M parameters), FractalMamba++ achieves 48.8 mIoU (SS) and 49.7 mIoU (MS), outperforming prior models such as Spatial Mamba (48.6/49.4), MSVMamba (47.6/48.5), and FractalMamba (48.0/48.9). At the Medium scale (44M), it achieves 45.8/46.2, exceeding MSVMamba-M (45.1/45.4) and Efficient VMamba (41.5/42.1). At the Nano scale (37M), it obtains 44.5/45.0, outperforming Local-ViM (43.4/44.4) and PlainMamba (44.1/–). These results demonstrate the strong segmentation capability of FractalMamba++ across different model sizes. The consistent performance gain over existing Mamba-based backbones highlights the benefits of its fractal-based serialization, long-range routing via CSR, and enhanced positional encoding through PRC.
@@ 4.5 Ablation Studies
We conduct ablation studies of our FractalMamba++ across various resolutions on the ImageNet-1K dataset. Classification performance is evaluated under multiple input resolutions: 2242, 3842, 6402, and 10242.
@@ 4.5.1 Effectiveness of Proposed Modules
As shown in Table VII, introducing the Fractal curve alone yields consistent improvements, particularly under higher resolutions (e.g., +9.4% at $1024^{2}$), indicating its effectiveness in maintaining spatial structure during serialization. Adding the CSR mechanism further enhances performance, especially at larger scales (e.g., +3.9% at $1024^{2}$), by facilitating long-range interactions and mitigating information fading. Incorporating the PRC module provides the final boost, bringing performance to 84.1% at $224^{2}$ and 78.8% at $1024^{2}$, highlighting its ability to reinforce relative positional cues and address residual structural disruptions caused by curve inflections. These results validate the complementary nature of the proposed modules and demonstrate their effectiveness in enhancing the resolution adaptability and spatial modeling capacity of FractalMamba++.
@@ 4.5.2 Effectiveness of Fractal Curves in Mamba Backbone
Table VIII shows that fractal serialization significantly improves classification performance, particularly under high-resolution inputs. For ViM, the top-1 accuracy at $1024^{2}$ increases from 16.1% to 39.7%, indicating that the fractal curve effectively preserves spatial continuity that would otherwise be lost in naive raster scanning. Similarly, for VMamba, accuracy improves from 62.9% to 72.3% at $1024^{2}$. These results confirm that fractal-based patch serialization enhances the multi-resolution robustness of both convolution-free and SSM-based architectures, demonstrating its generalizability and effectiveness across different model families.
@@ 4.5.3 Effectiveness of CSR in Mamba Backbone
TABLE X: Ablation study on different types of fractal curves in VMamba. Hilbert, Coil, and Meurthe curves are evaluated under varying input resolutions.
@@
Hilbert 83.0 83.3 83.0 78.8
Coil 82.8 83.1 82.6 78.5
Meurthe 82.9 83.0 82.8 78.5
@@ 4.5.4 Comparison of Fractal Curve Types
To assess the impact of different fractal serialization strategies, we replace the Hilbert curve in VMamba with two alternative fractal curves: Coil and Meurthe. As shown in Table X, Hilbert consistently outperforms the other two curves across all resolutions. At 10242, Hilbert achieves 78.8 accuracy, while Coil and Meurthe reach 78.3 and 78.1, respectively. Although the performance gap is small, the results suggest that Hilbert is slightly more effective in preserving spatial locality during serialization. Nonetheless, the competitive results of Coil and Meurthe demonstrate that our framework is robust to the choice of fractal curve.
@@
MS-Train 83.1 83.2 80.4 72.0
Fractal Only 83.0 83.3 83.0 78.8
@@ 4.5.5 Fractal-based Modeling versus Multi-scale Training
To evaluate whether fractal-based modeling can serve as a substitute for multi-scale training, we conduct experiments on ImageNet-1K using GrootV as the backbone. In the multi-scale training (MS-Train) setup, input resolutions are randomly sampled from ${224^{2},384^{2},512^{2}}$ during training. In contrast, the fractal-only setup trains on a fixed resolution of $224^{2}$ but applies fractal serialization to preserve spatial structure.
Table XI shows that both methods perform similarly at low resolutions. For example, at $224^{2}$ and $384^{2}$, the accuracy of fractal modeling is comparable to that of MS-Train. However, as input resolution increases, fractal-based modeling demonstrates significantly better generalization, achieving 83.0% at $640^{2}$ and 78.8% at $1024^{2}$, outperforming MS-Train by 2.6 and 6.8 points, respectively. These results suggest that fractal serialization can better maintain structural consistency across scales, offering a resolution-robust alternative to traditional multi-scale training.
@@ 5 Conclusion
This paper presents FractalMamba++, a novel vision backbone that enhances the Mamba architecture with improved adaptability to spatial structure and resolution variability. Motivated by the limitations of existing linear serialization strategies, we introduce a fractal-based patch scanning approach that better preserves local spatial relationships during 2D-to-1D conversion. To further address long-range dependency degradation and structural discontinuities introduced by fractal inflection points, we incorporate two dedicated modules: Cross-State Routing (CSR) for enhanced global information propagation and Positional-Relation Capture (PRC) for explicit modeling of spatial adjacency. Extensive experiments demonstrate the effectiveness and robustness of FractalMamba++ across a wide range of visual tasks and resolutions. Our method consistently outperforms existing Mamba-based backbones in multi-resolution image classification, achieving competitive performance at low resolutions while maintaining strong accuracy under extreme-resolution inputs.

### v2

@@ Abstract
Vision Mamba has recently emerged as a promising alternative to Transformer-based architectures, offering linear complexity in sequence length while maintaining strong modeling capacity. However, its adaptation to visual inputs is hindered by challenges in 2D-to-1D patch serialization and weak scalability across input resolutions. Existing serialization strategies such as raster scanning disrupt local spatial continuity and limit the model’s ability to generalize across scales. In this paper, we propose FractalMamba++, a robust vision backbone that leverages fractal-based patch serialization via Hilbert curves to preserve spatial locality and enable seamless resolution adaptability. To address long-range dependency fading in high-resolution inputs, we further introduce a Cross-State Routing (CSR) mechanism that enhances global context propagation through selective state reuse. Additionally, we propose a Positional-Relation Capture (PRC) module to recover local adjacency disrupted by curve inflection points. Extensive experiments across diverse downstream tasks, including image classification, semantic segmentation and object detection, demonstrate that FractalMamba++ consistently outperforms previous Mamba-based backbones, with particularly notable gains under high-resolution settings.
@@ 1 Introduction
Figure 1: Visualization of correlation between the final state and intermediate state tokens in the deepest feature layer under different input resolutions. The $16\times$ and $32\times$ maps correspond to the output token grids for input resolutions of $512\times 512$ and $1024\times 1024$, respectively. Without CSR, correlations are biased and localized due to information fading across long sequences. With CSR enabled, strong and distributed correlations are observed, demonstrating that global contextual information is effectively propagated across the sequence. “W/o” means without operation.
@@
Figure 2: Illustration of structural limitation of fractal curves. Certain corner points along the fractal traversal path (highlighted by purple and red color) experience disrupted adjacency relationships. Although spatially close in the 2D image grid, these patches become distant in the serialized sequence, impairing local continuity and information exchange.
@@
  • We conduct a more comprehensive set of experiments than prior work, evaluating FractalMamba++ across multiple parameter scales on diverse visual tasks including classification, segmentation, detection.
@@ 3.3 Cross-State Routing
TABLE II: Image classification performance (Top-1 Accuracy) on ImageNet-1k under varying input resolutions. FLOPs are measured at input resolution of $224\times 224$. Following the setting of MSVMamba [34], we evaluate three model capacities: Tiny (FractalMamba++ (T)), Micro (FractalMamba++ (M)), and Nano (FractalMamba++ (N))
Method Publication Param. FLOPs Input Resolution
2242 2562 3842 5122 6402 7682 10242 12802 14082 15362
Mamba® CVPR’25 29M 4.6G 81.1 45.7 25.4 12.8 7.8 5.3 2.8 1.8 1.6 1.4
FractalMamba++ (T) Year’25 30M 4.8G 83.0 83.5 84.1 83.9 83.0 81.9 78.8 74.3 71.3 67.5
FractalMamba++ (M) Year’25 11M 1.6G 79.5 80.6 82.0 81.3 80.1 78.3 73.3 66.3 61.7 56.1
ViM ICML’24 7M 1.5G 76.1 76.3 70.4 67.4 51.4 30.6 16.1 7.2 4.1 1.8
FractalMamba++ (N) Year’25 7M 1.0G 77.3 78.4 79.5 78.4 76.4 73.7 66.5 55.2 48.1 42.5
@@ 3.4 Positional-Relation Capture
where $\operatorname{Re}[\cdot]$ denotes the real part and is the complex conjugate. This formulation introduces a rotation-phase difference proportional to the relative distance (nm)(n-m), thereby encoding relative positions directly into token similarity.
@@ 4 Experiments
We conduct a series of experiments to evaluate and compare FractalMamba++ against several established Mamba-based benchmark models across a diverse range of vision tasks. These tasks include image classification, object detection, remote sensing binary change object detection, and semantic segmentation. Each task requires a different input resolution, which provides a comprehensive testbed to evaluate the multi-scale adaptability of our model. In particular, the remote sensing binary change object detection task involves high-resolution input images of size 1024 $\times$ 1024. This setting places strong demands on the model’s ability to maintain spatial consistency and capture long-range structural relationships. All experiments are conducted using 8 NVIDIA H800 GPUs. Following the setting of MSVMamba [34], we evaluate three model capacities: Tiny (FractalMamba++ (T)), Micro (FractalMamba++ (M)), and Nano (FractalMamba++ (N))
@@ 4.1 Image Classification
FractalMamba++ consistently achieves competitive or superior performance across all tested resolutions. At lower resolutions (e.g., $224^{2}$, $384^{2}$), it performs on par with state-of-the-art Mamba-based models. However, as input resolution increases, FractalMamba++ shows significantly slower degradation in accuracy compared to all baselines. For instance, at $1024^{2}$, the 30M variant of FractalMamba++ achieves 78.8 top-1 accuracy, outperforming SpatialMamba (66.1), GrootV (67.9), and the original VMamba (62.3). Similar trends are observed at 11M and 7M scales. This superior scalability highlights the effectiveness of our design. The fractal-based patch serialization preserves spatial locality across scales, while the CSR and PRC modules enhance long-range context propagation and mitigate structural disruptions caused by curve inflections. Together, these modules allow FractalMamba++ to maintain strong performance under extreme-resolution inputs, confirming its multi-scale adaptability and robustness for real-world visual applications.
To complement accuracy evaluations, we further compare the inference efficiency of various Mamba-based models on ImageNet-1K at $224\times 224$ resolution. As shown in Table III, we report frames per second (FPS) on a single H800 GPU with batch size 128. FractalMamba++ achieves competitive inference speeds within each parameter scale, while offering consistently superior multi-resolution performance. In the $\sim$30M group, FractalMamba++ runs at 302 FPS, slower than VMamba (1916 FPS) and MambaVision (3426 FPS), but on par with GrootV (315 FPS). In the medium-scale group ($\sim$11M), FractalMamba++ achieves 511 FPS, and outperforms ViM (398 FPS) in the small-scale group ($\sim$7M). Although FractalMamba++ incurs moderate latency due to additional modules (CSR, PRC), it still delivers a favorable trade-off between speed and accuracy, especially under high-resolution scenarios where other models degrade sharply. Notably, the current implementation of the CSR module involves custom CUDA operators, which, while functional, have not been fully optimized. This suggests that there remains considerable headroom for further acceleration through kernel-level refinement. This highlights the practical deployment potential of FractalMamba++ in resource-constrained settings.
@@ 4.2 Object Detection
Table IV presents detection and segmentation results for FractalMamba++ and several representative Mamba-based models, including VMamba [26], MSVMamba [34], Efficient VMamba [38], SpatialMamba [35], DefMamba [56], and FractalMamba [32]. Performance is reported in terms of mean average precision (mAP) for bounding boxes and instance masks. FractalMamba++ achieves consistent improvements under both training settings. Under the 1× schedule, the 41M model obtains 47.8 mAP for box prediction and 43.2 mAP for mask prediction, outperforming Spatial Mamba (47.6 and 42.9), MSVMamba (46.9 and 42.2), and VMamba (46.5 and 42.1). Under the 3× schedule, FractalMamba++ further improves to 49.5 for box mAP and 44.1 for mask mAP, achieving the highest performance among comparable models. Similar trends are observed at the 31M scale, where FractalMamba++ achieves 46.8 for box mAP and 42.2 for mask mAP, compared with 46.3 and 41.8 from MSVMamba. These results confirm the effectiveness and generalizability of FractalMamba++ across dense prediction tasks. The fractal-based serialization improves spatial structure retention, while the CSR and PRC modules enable stronger context modeling and structural continuity, which are essential for high-resolution visual understanding.
TABLE V: Semantic segmentation results on ADE20K using UPerNet with input resolution $512\times 2048$. FLOPs are computed under the same setting. $S$ and MS denote single-scale and multi-scale inference.
@@ 4.3 Semantic Segmentation
Table V reports results across three model capacities. At the tiny scale (62M parameters), FractalMamba++ achieves 48.8 mIoU (SS) and 49.7 mIoU (MS), outperforming prior models such as Spatial Mamba (48.6/49.4), MSVMamba (47.6/48.5), and FractalMamba (48.0/48.9). At the micro scale (44M), it achieves 45.8/46.2, exceeding MSVMamba-M (45.1/45.4). At the nano scale (37M), it obtains 44.5/45.0, outperforming Local-ViM (43.4/44.4) and PlainMamba (44.1/44.6). These results demonstrate the strong segmentation capability of FractalMamba++ across different model sizes. The consistent performance gain over existing Mamba-based backbones highlights the benefits of its fractal-based serialization, long-range routing via CSR, and enhanced positional encoding through PRC.
@@ 4.5 Ablation Studies
We conduct ablation studies of our FractalMamba++ across various resolutions on the ImageNet-1K dataset. Classification performance is evaluated under multiple input resolutions: 2242, 3842, 6402, and 10242. All experiments are conducted using the Tiny configuration (FractalMamba++ (T)).
@@ 4.5.1 Effectiveness of Proposed Modules
As shown in Table VII, introducing the Fractal curve alone yields consistent improvements, particularly under higher resolutions (e.g., +10% at $1024^{2}$), indicating its effectiveness in maintaining spatial structure during serialization. Adding the CSR mechanism further enhances performance, especially at larger scales (e.g., +3.9% at $1024^{2}$), by facilitating long-range interactions and mitigating information fading. Incorporating the PRC module provides the final boost, bringing performance to 84.1 at $384^{2}$ and 78.8 at $1024^{2}$, highlighting its ability to reinforce relative positional cues and address residual structural disruptions caused by curve inflections. These results validate the complementary nature of the proposed modules and demonstrate their effectiveness in enhancing the resolution adaptability and spatial modeling capacity of FractalMamba++.
@@ 4.5.2 Effectiveness of Fractal Curves in Mamba Backbone
Table VIII shows that fractal serialization significantly improves classification performance, particularly under high-resolution inputs. For ViM, the top-1 accuracy at $1024^{2}$ increases from 16.1 to 39.7, indicating that the fractal curve effectively preserves spatial continuity that would otherwise be lost in naive raster scanning. Similarly, for VMamba, accuracy improves from 62.3 to 72.3 at $1024^{2}$. These results confirm that fractal-based patch serialization enhances the multi-resolution robustness of both convolution-free and SSM-based architectures, demonstrating its generalizability and effectiveness across different model families.
@@ 4.5.3 Effectiveness of CSR in Mamba Backbone
TABLE X: Ablation study on different types of fractal curves in FractalMamba++. Hilbert, Coil, and Meurthe curves are evaluated under varying input resolutions.
@@
Hilbert 83.0 84.1 83.0 78.8
Coil 82.8 83.9 82.6 78.3
Meurthe 82.9 84.0 82.8 78.1
@@ 4.5.4 Comparison of Fractal Curve Types
To assess the impact of different fractal serialization strategies, we replace the Hilbert curve in FractalMamba++ with two alternative fractal curves: Coil and Meurthe. As shown in Table X, Hilbert consistently outperforms the other two curves across all resolutions. At 10242, Hilbert achieves 78.8 accuracy, while Coil and Meurthe reach 78.3 and 78.1, respectively. Although the performance gap is small, the results suggest that Hilbert is slightly more effective in preserving spatial locality during serialization. Nonetheless, the competitive results of Coil and Meurthe demonstrate that our framework is robust to the choice of fractal curve.
@@
MS-Train 83.0 83.1 79.5 62.4
Fractal Only 82.7 82.8 80.6 72.3
@@ 4.5.5 Fractal-based Modeling versus Multi-scale Training
To evaluate whether fractal-based modeling can serve as a substitute for multi-scale training, we conduct experiments on ImageNet-1K using VMamba as the backbone. In the multi-scale training (MS-Train) setup, input resolutions are randomly sampled from $224^{2},384^{2},512^{2}$} during training. In contrast, the fractal-only setup trains on a fixed resolution of $224^{2}$ but applies fractal serialization to preserve spatial structure.
Table XI shows that both methods perform similarly at low resolutions. For example, at $224^{2}$ and $384^{2}$, the accuracy of fractal modeling is comparable to that of MS-Train. However, as input resolution increases, fractal-based modeling demonstrates significantly better generalization, achieving 80.6 at $640^{2}$ and 72.3 at $1024^{2}$, outperforming MS-Train by 1.1 and 9.9 points, respectively. These results suggest that fractal serialization can better maintain structural consistency across scales, offering a resolution-robust alternative to traditional multi-scale training.
TABLE XII: Generalization performance on the downstream HRSOD task.
Models HRSOD
[…]
PGNet (Swin-T) 0.931 46.92
PGNet (FractalMamba++) 0.947 44.89
4.5.6 Generalization to Downstream Tasks
To further assess the generalization capability of FractalMamba++, we further conduct experiments on the downstream task: high-resolution salient object detection (HRSOD). For HRSOD, we employ the PGNet [60] framework, replacing its original Swin-T backbone with FractalMamba++, and evaluate performance using the maxF and BDE metrics. As shown in Table XII, FractalMamba++ demonstrates strong generalization across both tasks. On HRSOD, it achieves a maxF score of 0.947 and a BDE of 44.89, surpassing the Swin-T baseline (0.931 / 46.92). These results highlight the potential of FractalMamba++.
@@ 5 Conclusion
This paper presents FractalMamba++, a novel vision backbone that enhances the Mamba architecture with improved adaptability to spatial structure and resolution variability. Motivated by the limitations of existing linear serialization strategies, we introduce a fractal-based patch scanning approach that better preserves local spatial relationships during 2D-to-1D conversion. To further address long-range dependency degradation and structural discontinuities introduced by fractal inflection points, we incorporate two dedicated modules: CSR for enhanced global information propagation and PRC for explicit modeling of spatial adjacency. Extensive experiments demonstrate the effectiveness and robustness of FractalMamba++ across a wide range of visual tasks and resolutions. Our method consistently outperforms existing Mamba-based backbones in multi-resolution image classification, achieving competitive performance at low resolutions while maintaining strong accuracy under extreme-resolution inputs.